Home  Contact  Equipment  Field Notes  Photos  Data  Logs Today's Date:   
   
   
  Biomet public Home
Latest Entries
 TZ 2015-01-21 - SM
 SI 2015-01-20 - PO
 TW 2015-01-20 - PO
 VR 2014-12-09 - JV
 MB 2015-01-20 - PO
 WP 2015-01-20 - SK
 TA 2015-01-20 - PO
 EE 2015-01-20 - SK

Recently Added
 MB 2015-01-20 - PO
 SI 2015-01-20 - PO
 TA 2015-01-20 - PO
 TW 2015-01-20 - PO
 TZ 2015-01-21 - SM

   
   
   
   
   
 
 
<--2015-01-07 10:00:00 | -->

Tonzi Ranch: 2015-01-21 09:45:00
Author: Siyan Ma
Others: Joe

Summary: Updated CR1000 OS, Restarted co2 profile system, new CO2 tank, standard fieldwork, clip plots


When: Date: 1/21/2015 , Day: 21 , Watch Time: 9:45 am .
Where: Tonzi/Vaira Ranch
Who: Siyan, Joe
Weather: cloudy
Phenology:

image001.jpgimage002.jpgimage003.jpgimage004.jpg

Summary
  1. Joe updated the operating system of CR1000 and swapped memory cards of CR1000s at three sites.
  2. Checked and downloaded winflux and met data from PC at Tower, Floor, and Vaira.
  3. The CO2 profile system back to work after Joe fixed the analyzer in the lab.
  4. Read dendrometers
  5. Measured soil moisture
  6. Collect grass samples
  7. Measure groundwater well


Vaira_Grass Site
  1. Downloaded Phenocam images.
  2. Joe updated the operating system of CR1000 and swapped the memory cards of CR1000.
  3. The fans of air temp and relative humidity sensors did not working.
  4. Downloaded met and soil CO2 profile data:

image005.jpg

image006.jpg
  1. Checked levels of Radiation sensors, clear up surfaces.YES!
  2. Checked rain gage, make sure not clogged. YES! Joe put a screen for rain gage.
  3. Closed gate. YES!

Tonzi_Tower Site

  1. Fence and gate power: 12 v.
  2. Downloaded Winflux data from PC. YES!

image007.jpg
  1. swapped the memory cards of CR1000.

image008.jpg
  1. download Phenocam images. Yes!
  2. checked the fans of air temp. and relative humidity sensors on the tower and the floor. YES!
  3. Checked rain gage, make sure not clogged. YES!
  4. Downloaded met data from PC

image009.jpg
  1. The CO2 profile is back to work after Joe fixed the analyzer in the lab. New CO2 tank.

9. closed gate and set up fence. YES!
Toniz_Floor Site
  1. Fence and gate power: 12 v.
  2. downloaded Winflux data from PC. YES!

image010.jpg
  1. Swapped the memory cards of CR1000.

image011.jpg
  1. downloaded met data from PC. YES!

CR23X3 – floor-Met
CR23X6 – tower-Met
CR10X2 – floor-Met
Soil CO2 profile: CR23X5 – soil CO2 profile
image012.jpg

image013.jpg

image014.jpg

image015.jpg
  1. Closed gate and set up fence. YES!

Field Data

CO2 Profile Tanks

DateTankPressure (psi)ppm
2015-01-21CO2
2015-01-21N2 13500

Well Depth

DateLocationDepth (feet)
2015-01-21ROAD38.5
2015-01-21TOWER34.7
2015-01-21FLOOR31.3

Moisture Point Probes TZ
LocationRep0-1515-3030-4545-60
Floor_South114.128.637.750
Floor_South214.128.937.448.8
Floor_SouthAVG14.128.837.649.4
Floor_West122.631.337.446
Floor_West2223138.246.3
Floor_WestAVG22.331.237.846.2
SoilCO2129.729.241.340
SoilCO2229.430.140.739.7
SoilCO2AVG29.629.74139.9
Soil_Transect_4120.9242428.8
Soil_Transect_4222.624.323.529.4
Soil_Transect_4AVG21.824.223.829.1
Soil_Transect_3123.530.127.637.4
Soil_Transect_3224.730.427.638
Soil_Transect_3AVG24.130.327.637.7
Soil_Transect_2116.830.13146.6
Soil_Transect_2216.530.132.146.8
Soil_Transect_2AVG16.730.131.646.7
Soil_Transect_1118.218.231.827.6
Soil_Transect_1218.518.532.428.2
Soil_Transect_1AVG18.418.432.127.9
SoilCO2_West_A123.825.532.741.4
SoilCO2_West_A224.725.214.342.3
SoilCO2_West_AAVG24.325.423.541.9
SoilCO2_West_B123.816.131.847.7
SoilCO2_West_B224.716.43148
SoilCO2_West_BAVG24.316.331.447.9
SoilCO2_West_C121.228.933.245.4
SoilCO2_West_C221.729.834.647.1
SoilCO2_West_CAVG21.529.433.946.3
Floor_East123.639.641.4
Floor_East229.835.7
Floor_EastAVG46.326.737.741.4
Moisture Point Probes VR
LocationRep0-1515-3030-4545-60
Solar_Panels113.829.827.635.7
Solar_Panels215.327.128.539.4
Solar_PanelsAVG14.628.528.137.6
Flux_Tower11.926.120.927.4
Flux_Tower218.220.726.2
Flux_TowerAVG10.126.120.826.8
Gate113.515.520.715.6
Gate213.515.819.816.2
GateAVG13.515.720.315.9
Biomass_Cage16.817.927.927.6
Biomass_Cage26.817.929.928.2
Biomass_CageAVG6.817.928.927.9

No VR grass heigth found

No TZ grass heigth found

Tree DBH

DateTag#Reading (mm)
2015-01-21111.5
2015-01-211027.7
2015-01-21213.0
2015-01-2124519.6
2015-01-21313.2
2015-01-2149.0
2015-01-21512.9
2015-01-21710.0
2015-01-217219.5
2015-01-217915.1
2015-01-21813.5

Pending

No TZ Grass Biomass data found

No TZ OakLeaves Biomass data found

No TZ PineNeedles Biomass data found

0 photos found

Graphs display: 7 days | 14 days | 28 days
10 graphs found

No data found for TZ_Fenergy.
8 TZ_tmet graphs found

11 TZ_soilco2 graphs found

Can't check TZ_mixed
2 TZ_mixed graphs found

1 TZ_sm graphs found

1 TZ_fflux graphs found

5 TZ_fmet graphs found

1 TZ_tram graphs found

No data found for TZ_co2prof.
Can't check badm_TZ
2 badm_TZ graphs found

Tonzi Z-cam data


 
 

Login
Select Notes:

Search Title for:

Search Descritption for:

This material is based upon work supported by the National Science Foundation and US Department of Energy. Any opinions, findings, conclusions, or recommendations expressed in the material are those of the author(s) and do not necessarily reflect the views of the supporters.